OpenStudy (anonymous):

If the brakes of an automobile are applied from 70 km/h to 35 km/h, how many meters does the car travel during this deceleration?

OpenStudy (anonymous):

Initial vel u = 19.40 m/s (70km/h) Final vel v = 9.70 m/s Average vel = ½(u +v) = d/t d = ½(u+v)t .. .. ½(29.10) x t .. .. ►d = 14.55t .. where t = time in sec...!!HOPE YOU GET THIS.